OpenStudy (anonymous):

Find K such that x^2 +K+4=0 shall be: 1. Real and Equal 2. Real and Different 3.Complex

sam (.sam.):

Real and Equal

OpenStudy (anonymous):

use the discriminant ... do u know how?

OpenStudy (anonymous):

REAL AND EQUAL D=0

OpenStudy (anonymous):

I do, thank you.

OpenStudy (anonymous):

okee dokee

OpenStudy (anonymous):

I have just reached a point of uncertainty. Could you assist me a bit?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Could you guide me through coming to my Real and equal answer? I used the discriminant coming up with \[(-2)^2-4(K)(3)\] What am I missing here?

OpenStudy (anonymous):

D=b^2-4ac

OpenStudy (anonymous):

b is 0 - no x term

OpenStudy (anonymous):

Well if this is correct, I have K=32. Somewhere around the ballpark?

OpenStudy (anonymous):

D= 0 - 4(1)(k+4)=0 and solve for k

OpenStudy (anonymous):

-4k -16 =0 , k =-4

OpenStudy (anonymous):

so... Discriminant is the quantity under the √ in the quadratic formula - b^2-4ac, in your case a = 1 , b = 0 , c = K+4

OpenStudy (anonymous):

real and equal D=0, real and different D>0, complex D<0

OpenStudy (anonymous):

How would I know that I need to put it in that format though?

OpenStudy (anonymous):

the discriminant determines the nature of the roots

OpenStudy (anonymous):

[i.e. D=0-4(1)(k+4)] I'm not really sure how you came about it really. The way my professor taught it, he just gave us the natures and what you just said: real and equal D=0, real and different D>0, complex D<0

OpenStudy (anonymous):

so what is D?

OpenStudy (anonymous):

x^2 +0x+(K+4)=0 ---- this is the complete equation

OpenStudy (anonymous):

D is the discriminant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is equal to b^2-4ac ... but your b is 0

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

Yes I am starting to. So in 0-4(1)(k+4)=0 What made you use k+4 in a parenthesis?

OpenStudy (anonymous):

so I can make it obvious that K+4 is c

OpenStudy (anonymous):

did you understand now?